# ChainRank-DPO: Chain Rank Direct Preference Optimization for LLM Rankers

### v1

@@ Abstract
Large language models (LLMs) have demonstrated remarkable effectiveness in text reranking through works like RankGPT, leveraging their human-like reasoning about relevance. However, supervised fine-tuning for ranking often diminishes these models’ general-purpose capabilities, including the crucial reasoning abilities that make them valuable for ranking. We introduce a novel approach integrating Chain-of-Thought prompting with an SFT-DPO (Supervised Fine-Tuning followed by Direct Preference Optimization) pipeline to preserve these capabilities while improving ranking performance. Our experiments on TREC 2019 and 2020 Deep Learning datasets show that our approach outperforms the state-of-the-art RankZephyr while maintaining strong performance on the Massive Multitask Language Understanding (MMLU) benchmark, demonstrating effective preservation of general-purpose capabilities through thoughtful fine-tuning strategies. Our code and data will be publicly released upon the acceptance of the paper.
ChainRank-DPO: Chain Rank Direct Preference Optimization for
LLM Rankers
@@ 1 Introduction
Text reranking is a vital task in information retrieval Liu (2009); Hasanain (2018), crucial for search engines Li et al. (2024), conversational AI Becker et al. (2012), and recommendation systems Zhao and Liu (2024). Large language models (LLMs) excel in reranking due to their reasoning and human-like thinking capabilities, enabling them to handle complex queries and ambiguous contexts. RankGPT Achiam et al. (2023) has set a high standard in listwise reranking Ma et al. (2019), leveraging its interpretability and systematic reasoning Brown et al. (2020) to advance reranking technologies.
Figure 1: While both methods can rerank passages, general methods struggle with simple QA tasks, whereas ChainRank successfully solves them step by step.
@@
Building on the success of RankGPT, several models have sought to distill its output into smaller, task-optimized models through supervised fine-tuning (SFT). RankVicuna Pradeep et al. (2023a), for instance, uses distilled data generated exclusively from GPT-3.5, focusing on efficient fine-tuning to improve ranking performance while maintaining a lightweight architecture. Moreover, RankZephyr Pradeep et al. (2023b) employs a more comprehensive approach utilizing distilled data from GPT-3.5 and GPT-4, ensuring the quality and reliability of training data to achieve superior ranking results. Although these models achieve state-of-the-art performance in specific benchmarks, they highlight a key challenge in ranking: the trade-off between optimizing task-specific performance and preserving the broader reasoning capabilities of LLMs. Addressing this challenge requires novel strategies to balance specialization with generalization, ensuring that ranking models retain their versatility across diverse tasks.
Our investigations in Figure 1 reveal that while RankVicuna and RankZephyr exhibit strong performance in listwise ranking tasks, they have forfeited their mathematical reasoning capabilities following fine-tuning. This is evident from the figure, where the models fail to produce coherent solutions to mathematical problems, often outputting unrelated or nonsensical numerical sequences. Moreover, further analysis uncovered that these models have also lost their general text-generating capabilities. Regardless of the input prompt, the models struggle to produce meaningful or contextually appropriate responses, highlighting a significant trade-off introduced by their specialized fine-tuning for ranking tasks. Furthermore, while they exhibit comprehension of listwise ranking, they encountered difficulties with pairwise ranking, as their pairwise outcomes did not correspond with those generated by the listwise ranking Peng et al. (2024).
To address this limitation, we propose ChainRank strategy, built on LLaMA3-8b-instruct Dubey et al. (2024). Using the same training data as RankZephyr and RankGPT3.5/RankGPT4 as teacher models, we introduced a chain-of-thought (CoT) prompt to guide sequential passage ranking. We then implemented ChainRank-DPO, enhancing the model’s reasoning abilities for superior ranking performance. While traditional Direct Preference Optimization (DPO) Rafailov et al. (2024) performs well on chat benchmarks, it struggles with long-chain reasoning tasks like math and ranking due to error propagation. Inspired by Step-DPO Lai et al. (2024), we designed a novel DPO framework using overlapping ranking orders as the reward function, allowing better error correction. ChainRank preserves general text generation and reasoning capabilities, overcoming limitations in RankVicuna and RankZephyr. It outperforms both on TREC 2019/2020 Deep Learning Tracks Craswell et al. (2021, 2020) and BEIR benchmarks Thakur et al. (2021). We summarize the contributions of this work as follows:
@@
  • We propose a novel chain-of-thoughts instruction (reranking) tuning that enables LLMs to rank passages based on relevance step by step.
  • To our best knowledge, this is the first work to introduce the Supervised Fine-Tune–DPO pipeline in the context of text reranking.
  • Using the same training data, our model outperforms both RankZephyr and the teacher model, RankGPT4, while retaining the general capabilities of LLMs.
@@
Figure 2: ChainRank Chain-of-Thought (CoT) reranking prompt guiding the model to rank passages based on relevance to a query iteratively. The prompt ensures step-by-step selection, removal, and ordering of passages, with an example illustrating the expected output format.
@@ 2 Methodology
Figure 3: Training framework of our model. $Left:$ Stage 1 aims to teach the student model to perform text reranking. $Right:$ In stage 2, we use the same prompt format to generate multiple answers, then pick the chosen and rejected answers based on the number of overlapped steps, finally, we utilize the preference data to perform Chain DPO training.
2.1 CoT Reranking Prompt Design
Recent advancements in listwise text reranking, such as RankVicuna Pradeep et al. (2023a), RankZephyr Pradeep et al. (2023b), and related methods Liu et al. (2024), build on zero-shot techniques Ma et al. (2023); Sun et al. (2023), using prompt templates to reorder documents for metrics like nDCG Järvelin and Kekäläinen (2002). Given a query $q$ and documents ${d_{1},d_{2},...,d_{n}}$, the goal is to return a reordered list (e.g., $[1]>[2]>[3]...>[20]$).
Our ChainRank strategy frames listwise reranking as a chain-of-thought (CoT) reasoning task, selecting the most relevant document iteratively until all are ranked. The CoT prompt for LLaMA is shown in Figure 2.
2.2 Training Dataset
We train ChainRank model using 35k GPT-3.5 and 5k GPT-4 labeled instances from Pradeep et al. (2023b), derived from randomly selected MS MARCO v1 queries. Pyserini Lin et al. (2021) retrieved 20 BM25 candidates per query, which RankGPT3.5 and RankGPT4 ordered as teacher models. Malformed generations, such as missing or duplicate identifiers, were removed to improve data quality. We randomly selected 90% of the query-document pairs for Stage 1 SFT training and 10% for Stage 2 DPO training, denoted as SFT data and DPO data.
2.3 Training Stage 1: Supervised Fine-Tune
@@
As shown in Figure 3, during the Supervised Fine-Tuning (SFT) phase, we use SFT data, maintaining a zero-shot setup since RankGPT and ChainRank do not rely on human-labeled data.
@@
Before fine-tuning, the original LLaMA3-8B-Instruct model Dubey et al. (2024) failed to produce meaningful CoT reranking, simply replicating the example format without relevance-based results, highlighting the need for fine-tuning. The model weights are publicly available on HuggingFace. [^1] https://huggingface.co/meta-LLaMA/Meta-LLaMA-3-8B-Instruct
@@
We fully fine-tune the 8B parameter LLaMA-3 model for three epochs with a batch size of 128, a $5\times 10^{-6}$ learning rate, and bfloat16 format. Training on four NVIDIA A100 80GB GPUs took approximately 39 hours.
2.4 Training Stage 2: Chain DPO
@@
After Stage 1, the ChainRank-SFT model generates three ranking predictions ($y$) on prompts ($x$) from DPO training data. Predictions are evaluated by overlapping ranking orders with ground-truth labels, creating a preference dataset with prompts (x), chosen steps ($s_{w}$), rejected steps ($s_{l}$), and overlapping steps ($s_{o}$). Unlike prior methods, the final ranking $y$ comprises a sequence of reasoning steps, $y={s_{1},s_{2},\dots,s_{n}}$, where each step is conditioned on prior steps $\pi(s_{k}|x;s_{1:k-1})$. Overlapping steps ($s_{o}$) are tracked until a divergence is found, after which only the initial contiguous overlaps are included in $s_{o}$, and later steps are categorized as chosen ($s_{w}$) or rejected ($s_{l}$). This results in a dataset of $(x,s_{w},s_{l},s_{o})$.
@@
where $\pi_{\theta}(s\mid x;s_{o})$ represents the model’s probabilities, optimized to favor correct steps while leveraging reference model probabilities ($\pi_{\text{ref}}$) for stable training and better generalization, especially in zero-shot scenarios. This improves CoT reasoning’s sequential decision-making and model consistency. The Chain DPO stage is trained on four NVIDIA A100 80GB GPUs over one epoch, taking approximately eight hours.
@@ 3 Experiments
3.1 Research Questions
Before conducting the experiments, we formulate several research questions to ensure that ChainRank could effectively address the challenges:
-RQ1: Does the CoT instruction tuning improve the step-by-step relevance ranking of passages compared to traditional ranking methods?
-RQ2: How does our model’s performance on text reranking tasks compare with existing models like RankZephyr and RankGPT4 across various datasets and task settings?
-RQ3: Does introducing DPO improve the performance and robustness of our model?
@@ 3.2 Evaluation Benchmarks
We evaluate ranking capabilities using TREC DL19 Craswell et al. (2020) and DL20 Craswell et al. (2021) Tracks, derived from MS MARCO V1 Bajaj et al. (2016), with human-annotated relevance labels. TREC DL19 and DL20 contain 43 and 54 queries, respectively, each paired with 100 candidate passages retrieved by BM25. Additionally, we include cross-domain datasets from BEIR Thakur et al. (2021), not seen by the fine-tuned LLMs, for broader analysis. Using a sliding window strategy with a size of 20 and a stride of 10, we rerank 100 passages per query.
We use the Massive Multitask Language Understanding (MMLU) benchmark Hendrycks et al. (2020) to evaluate whether our model retains its text understanding abilities after fine-tuning for ranking tasks. MMLU tests proficiency across 57 diverse subjects, from elementary to advanced professional topics. We compare our model’s performance on MMLU with baseline LLMs to ensure ranking fine-tuning does not compromise general-purpose language understanding and generation capabilities, maintaining its versatility in broader tasks.
@@ 3.3 Results
Table 1 compares model performance on TREC, BEIR, and MMLU benchmarks. Our 8B ChainRank model outperforms baselines, including RankGPT4  [^2] RankGPT is the only closed-source and largest model in the table., even with 90% of the training data, and surpasses RankLLaMA3 under identical training conditions, showcasing the effectiveness of our CoT reranking prompt. In the DPO stage, ChainRank-DPO improves across all datasets with just one training epoch, enhancing robustness to passage variations (Appendix B).
On MMLU, ChainRank matches LLaMA3’s performance, while RankLLaMA3 shows a slight drop and RankVicuna performs poorly. RankZephyr has completely lost its ability to generate meaningful outputs, it receives a score of 0 on MMLU. Additional examples are in Appendix D.
We present a figure to illustrate the trade-off between performance and inference cost when generating ranking orders at varying step intervals. Notably, our method achieves comparable latency to RankLLaMA3 (26.5 vs. 25.0 seconds). By utilizing parallel distributed evaluation, we further reduce inference latency to 6.73 seconds on 4 A100 GPUs, demonstrating the efficiency of our approach.
Models TREC BEIR MMLU DL19 DL20 NFC COVID FIQA AVG. RankGPT4 0.746 0.708 0.406 0.749 0.333 0.864 BM25 0.506 0.480 0.325 0.595 0.236 N/A Contriever 0.616 0.599 0.328 0.596 0.329 N/A Gemma-7B 0.533 0.530 0.338 0.573 0.236 0.649 LLaMA3 0.641 0.621 0.269 0.628 0.216 0.662 RankZephyr 0.742 0.709 0.331 0.592 0.230 0.000 RankVicuna 0.668 0.655 0.338 0.592 0.236 0.373 RankLLaMA3 0.730 0.635 0.335 0.672 0.330 0.628 ChainRank-SFT 0.752 0.714 0.353 0.766 0.342 0.663 ChainRank-DPO 0.755 0.717 0.358 0.772 0.342 0.663
Table 1: Performance of different models on TREC (nDCG@10) and MMLU (exact match score) benchmarks. All the reranking tasks are based on BM25 retrieval results. The bold values in the table highlight the best performance across the respective benchmarks.
Figure 4: FLOPs and nDCG@10 performance across different step intervals. The blue line represents the inference cost (FLOPs), while the orange line shows the ranking performance (NDCG@10). As the step interval increases, FLOPs decrease significantly, with a slight drop in nDCG@10 scores.
@@ 4 Conclusion and Future Work
In this paper, we propose a novel approach, ChainRank, a novel zero-shot listwise text reranking model built on LLaMA3. We demonstrated the effectiveness of CoT reranking prompt and our SFT–Chain DPO pipeline. The results of our experiments show that our model achieves superior performance compared to many open-source and closed-source LLMs, such as RankZephyr and RankGPT4.
In future work, we plan to explore the application of our method to other models, such as Mistral, Zephyr, and LLaMA3.1, to evaluate its generalizability and performance across different architectures. Additionally, our future research could incorporate higher-quality datasets with varying numbers of passages per instance to enhance the diversity and robustness of the training data. Expanding the dataset in this manner could lead to improved overall performance.

### v2

# RaCT: Ranking-aware Chain-of-Thought Optimization for LLMs

@@ Abstract
Large language models (LLMs) have shown significant promise in text reranking tasks by leveraging their advanced language understanding and reasoning capabilities. However, traditional supervised fine-tuning (SFT) approaches by ranking utilities can compromise LLMs’ general-purpose abilities. To address this challenge, we propose a novel LLM-based reranking algorithm – RaCT – that implements SFT with Chain-of-Thought prompting, followed by a ranking preference optimization (RPO). The proposed RaCT aims to enhance ranking performance for LLMs while preserving their inherent language modeling abilities. Experimental evaluations on the three public ranking benchmarks (TREC DL, BEIR, and BRIGHT) and one LLM benchmark demonstrate the superior ranking performance of RaCT with a retained language understanding and reasoning capacity.
RaCT: Ranking-aware Chain-of-Thought Optimization for LLMs
@@ 1 Introduction
Text reranking is a vital task in information retrieval Liu (2009); Hasanain (2018), crucial for search engines Li et al. (2024), conversational AI Becker et al. (2012), and recommendation systems Zhao and Liu (2024). Large language models (LLMs) excel in reranking due to their reasoning and human-like thinking capabilities, enabling them to handle complex queries and ambiguous contexts. RankGPT Achiam et al. (2023) has set a high standard in listwise reranking Ma et al. (2019), leveraging its systematic reasoning Brown et al. (2020) to advance reranking technologies.
Figure 1: General text reranking LLMs excel at ranking tasks but struggle with general tasks such as open-ended text generation. General LLMs (e.g., LLaMA) perform well on diverse tasks but lack strong ranking capabilities. RaCT effectively integrates both, achieving high performance in ranking while maintaining general language generation abilities.
@@
Building on top of RankGPT, several models have sought to distill its output into smaller, task-optimized models through supervised fine-tuning (SFT). To name a few, RankVicuna Pradeep et al. (2023a) uses distilled data generated exclusively from GPT-3.5 and focuses on efficient fine-tuning to improve ranking performance with a lightweight architecture. RankZephyr Pradeep et al. (2023b) develops a more comprehensive approach that utilizes distilled data from GPT-3.5 and GPT-4, ensuring the quality and reliability of training data to achieve superior ranking results. While these models exhibit state-of-the-art ranking performance, they pose a new challenge for LLM rerankers: the trade-off between optimizing ranking utility and preserving the inherent language modeling capabilities. Empirically, over-emphasizing ranking through SFT could significantly degrade LLM in general-purpose language understanding tasks, hindering its practical deployment as a reranking agent in real-world scenarios.
General rerankers achieve strong performance on listwise ranking tasks after SFT, but often lose their ability to follow general instructions. As shown in Figure 1, models like LLaMA may output incomplete or irrelevant results, such as numeric lists, when asked to write a travel blog. This suggests that SFT for ranking can reduce a model’s general text generation ability.
In this study, we introduce a novel LLM reranking algorithm, namely RaCT, by implementing SFT through a chain-of-thought (CoT) prompt (see Fig. 2) to guide sequential passage ranking without undermining the LLM’s general language understanding ability. To enhance the reasoning ability, we further incorporate CoT prompting into ranking preference optimization (RPO) Rafailov et al. (2024), proposing a new step-wise ranking preference learning framework RaCT. The framework utilizes overlapping ranking orders as the reward function, enabling efficient and flexible error correction.
We summarize the contributions of this work as follows:
  • To the best of our knowledge, this is the first research study to investigate the trade-off between ranking utility and language modeling for the recent emerging LLM re-rankers.
@@
  • We propose a novel chain-of-thoughts instruction (reranking) tuning algorithm – RaCT – that enables LLMs to rank passages based on relevance step by step, and RPO to enable further ranking preference optimization.
  • Empirical evidence on three ranking benchmarks, TREC Deep Learning Tracks Craswell et al. (2021a, 2020), BEIR Thakur et al. (2021), and BRIGHT Su et al. (2025), and the Massive Multitask Language Understanding (MMLU) benchmark Hendrycks et al. (2020) suggests that our approach achieves state-of-the-art LLM re-ranking performance with the preservation of intrinsic language modeling capabilities.
@@
Figure 2: RaCT Chain-of-Thought (CoT) reranking prompt guiding the model to rank passages based on relevance to a query iteratively. The prompt ensures step-by-step selection, removal, and ordering of passages, with an example illustrating the expected output format.
@@ 2 Methodology
Figure 3: Training framework of our model. $Left:$ Stage 1 aims to teach the student model to perform text reranking. $Right:$ In stage 2, we use the same prompt format to generate multiple answers, then pick the chosen and rejected answers based on the number of overlapped steps, finally, we utilize the preference data to perform RPO training.
2.1 CoT Reranking Prompt and Dataset
RaCT formulates listwise reranking as a chain-of-thought (CoT) reasoning task, where the model iteratively selects the most relevant passage until all are ranked. The prompt (Figure 2) guides LLaMA to output ranked indices given a query and a set of passages. This builds on prior zero-shot reranking methods such as RankVicuna and RankZephyr Pradeep et al. (2023a, b); Liu et al. (2024), which use template-based prompts to optimize ranking metrics like nDCG Järvelin and Kekäläinen (2002).
We train RaCT on 40k examples from Pradeep et al. (2023b), with 20 BM25-retrieved passages per query labeled by RankGPT3.5 or RankGPT4. The dataset includes three prompt formats: standard RankGPT-style prompts,RaCT CoT prompts, and RaCT CoT prompts where the model outputs the ranked list directly without explanation. We split 90% for supervised CoT tuning and 10% for RPO. Denoted as CoT tuning data and RPO data. The dataset is publicly available. [^1] https://huggingface.co/datasets/castorini/rank_zephyr_training_data
2.2 RaCT: Two-Stage Training
@@
Training Stage 1: CoT Tuning As shown in Figure 3, during the Stage 1, we use CoT Tuning data with mixed prompt templates, maintaining a zero-shot setup since RankGPT and RaCT do not rely on human-labeled data.
@@
Before fine-tuning, the LLaMA3.1-8B-Instruct model Grattafiori et al. (2024) failed to perform CoT reranking, merely copying the example format without relevance. This highlights the need for fine-tuning. Model weights are available on Hugging Face. [^2] https://huggingface.co/meta-LLaMA/Meta-LLaMA-3-8B-Instruct
@@
We fully fine-tune the 8B parameter LLaMA3.1 model for three epochs with a batch size of 128, a $5\times 10^{-6}$ learning rate, and bfloat16 format. Training on four NVIDIA A100 80GB GPUs took approximately 20 hours. We use Maximum Likelihood Estimation (MLE) as our objective function, optimizing the model to generate the most probable ranking sequence given the input passages and queries.
@@
Training Stage 2: RPO After Stage 1, the RaCT model generates three ranking predictions ($y$) on prompts ($x$) from RPO data. Predictions are evaluated by overlapping ranking orders with ground-truth labels, creating a preference dataset with prompts (x), chosen steps ($s_{w}$), rejected steps ($s_{l}$), and overlapping steps ($s_{o}$). Unlike prior methods, the final ranking $y$ comprises a sequence of reasoning steps, $y=\{s_{1},s_{2},\dots,s_{n}\}$, where each step is conditioned on prior steps $\pi(s_{k}|x;s_{1:k-1})$. Overlapping steps ($s_{o}$) are tracked until a divergence is found, after which only the initial contiguous overlaps are included in $s_{o}$, and later steps are categorized as chosen ($s_{w}$) or rejected ($s_{l}$). This results in a dataset of $(x,s_{w},s_{l},s_{o})$.
@@
where $\pi_{\theta}(s\mid x;s_{o})$ represents the model’s probabilities, optimized to favor correct steps while leveraging reference model probabilities ($\pi_{\text{ref}}$) for stable training and better generalization, especially in zero-shot scenarios. This improves CoT reasoning’s sequential decision-making and model consistency. The RPO stage is trained on four NVIDIA A100 80GB GPUs over one epoch, taking approximately 6 hours.
@@ 3.1 Evaluation Benchmarks
We evaluate ranking performance on standard academic benchmarks, including TREC DL19 and DL20 Craswell et al. (2020, 2021a), each with 100 BM25-retrieved passages per query and human relevance labels. To assess generalization, we include additional cross-domain benchmarks: BEIR Thakur et al. (2021), TREC DL21/DL22 Craswell et al. (2021b, 2022), and BRIGHT Su et al. (2025). Reranking is conducted using a sliding window strategy Sun et al. (2023), details explained in Appendix C. To verify general-purpose capabilities, we also evaluate on MMLU Hendrycks et al. (2020), testing 57 subject areas for text understanding post fine-tuning.
@@ 3.2 Results
Table 1 compares model performance on TREC and MMLU. Our 8B RaCT model outperforms all baselines, including the larger, closed-source RankGPT4333RankGPT is the only closed-source and largest model in the table., despite using only 90% of the training data. It also surpasses LLaMA3.1 + SFT, confirming the benefit of our CoT reranking prompt. In the RPO stage, RaCT achieves further gains with just one epoch and shows strong robustness: performance drops only 0.004 under tighter sliding windows, compared to over 0.02 for others.
We also applied LoRA to fine-tune only the linear layers in Stage 1, reducing training time by 4 hours while achieving a competitive 0.747 on DL19. This surpasses RankZephyr and RankGPT, showing that a lightweight fine-tuning approach can still achieve strong ranking performance with our CoT prompting strategy.
Models MSv1 MSv2 MMLU DL19 DL20 DL21 DL22 AVG. RankGPT4 0.750 0.704 0.684 0.509 0.864 BM25 0.506 0.480 0.446 0.269 N/A Gemma-7B 0.533 0.530 0.446 0.269 0.649 LLaMA3.1 0.669 0.641 0.639 0.429 0.720 DeepSeek-R1-8B 0.565 0.503 0.558 0.380 0.530 RankZephyr 0.742 0.709 0.600 0.408 0.000 RankVicuna 0.668 0.655 0.608 0.421 0.373 LLaMA3.1 + SFT 0.707 0.635 0.660 0.441 0.628 RaCT (Ours) 0.758 0.720 0.706 0.532 0.720
Table 1: Performance of different models on TREC (nDCG@10) and MMLU (exact match score) benchmarks. All the reranking tasks are based on BM25 retrieval results. The bold values highlight the best performance.
Model BEIR BRIGHT
BM25 0.426 0.137
RankZephyr 0.422 0.130
RaCT 0.545 0.178
Table 2: Average performance comparison on BEIR and BRIGHT benchmarks (nDCG@10), see detailed scores in Appendix F and G
Table 2 presents the performance comparison of RaCT against BM25 and RankZephyr across various BEIR and BRIGHt benchmark datasets. Both LLM models are reranking based on BM25 retrieval results. The results show that RankZephyr, despite being a fine-tuned model, receives a lower average score than the traditional retriever BM25, indicating that its reranking effectiveness does not always surpass basic retrieval methods. Combined with the results from DL21 and DL22, these findings suggest that while traditional fine-tuning methods improve performance on in-scope datasets, they do not necessarily generalize well to out-of-scope datasets. In contrast, RaCT demonstrates stronger generalization across different retrieval tasks, reinforcing the effectiveness of incorporating CoT prompting and RPO in improving reranking capabilities beyond the training distribution.
On MMLU, RaCT matches LLaMA3.1’s performance, while LLaMA3.1 w/o CoT shows a slight drop and RankVicuna performs poorly. RankZephyr has completely lost its ability to generate meaningful outputs, it receives a score of 0 on MMLU.
Although our model was trained using a mixed Chain-of-Thought (CoT) reranking prompt, we use the same concise prompt as RankGPT during inference and directly output the final ranked list. This ensures that inference time remains unchanged.
3.3 Ablation Study
Setting nDCG@10 (DL22)
[…]
– CoT 0.407
– RPO 0.504
– CoT&RPO 0.429
[…]
Table 3: Ablation study with LLaMA3.1-8B as the base model. Each row removes or modifies components from the full RaCT setup and reports performance on TREC DL22 (nDCG@10).
Table 3 presents an ablation study based on LLaMA3.1-8B. Removing CoT or RPO from the RaCTBM25 baseline leads to performance drops, with CoT having a larger impact (0.407 vs. 0.504). Removing both results in 0.429, which is lower than using CoT alone but higher than using RPO alone, indicating that both components contribute and their combination is most effective. Using SPLADE++ED as the retriever further boosts performance to 0.683, showing the reranker’s compatibility with stronger retrieval systems.
@@ 4 Conclusion and Future Work
In this paper, we introduced RaCT, a zero-shot listwise reranker based on LLaMA3.1, and showed the effectiveness of CoT prompting and CoT–RPO training. Our model outperforms both open- and closed-source baselines, including RankZephyr and RankGPT4. For future work, we plan to adapt RaCT to other architectures (e.g., Mistral, Zephyr, QWEN) and train on more diverse, higher-quality datasets to enhance robustness and performance.